OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5.

OpenStudy (saifoo.khan):

You have 2 cordinates, (6,0) and (0,5) You can solve it now.

OpenStudy (anonymous):

y=-5/6x+5

OpenStudy (anonymous):

slope-intercept form is y=mx+b, where m is the slope and b is the y-intercept, so you can just plug in the x-intercept as a point and plug in b. 0=m*6+5 -5=6m -5/6=m Then, plug the slope back into slope-intercept form to get... y=-5/6x+5